OpenStudy (anonymous):

Vectors: The guy wire supporting a 100 foot tower has a tension of 550 pounds. Using the distances: -50 and 75, write the component form of the vector F representing the tension in the wire.

sam (.sam.):

There's a diagram for this?

OpenStudy (anonymous):

Yes, there's a 3-D graph with the tower on the z-axis standing at 100 ft. The end of the guy wire has coordinates (75,-50,0)

OpenStudy (anonymous):

T_x=T*x/(SQRT(x^2+y^2+z^2)) T_y=T*y/(SQRT(x^2+y^2+z^2)) T_z=T*z/(SQRT(x^2+y^2+z^2))

OpenStudy (anonymous):

T_x=204.3 T_y=306.4 T_z=408.5

OpenStudy (anonymous):

I did not use plus or minus to signify directions.

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

I think so. I'm writing everything down now to make sure it's all working. Thank you!

OpenStudy (anonymous):

Your welcome. For vector components in 3 dimensions, think of the component you want to find and consider the rest one orthogonal component. That reduces it down to 2 dimensional angles.

OpenStudy (anonymous):

Oh, okay. That makes quite a bit more sense. The only thing I was thinking to do was to find the vector from (75,-50,0) to (0,0,100), which I knew wasn't right.